# Missing Data Imputation by Reducing Mutual Information with Rectified Flows

### v2

@@ 3.1 Mutual Information Reducing Iterations (MIRI)
Now we study the problem of training an imputer $\mathbf{g}$ to reduce the dependency between the imputed sample $\mathbf{X}(\mathbf{g})$ and $\mathbf{M}$. The approach adopted by (Liu et al., 2024) is to minimize the mutual information between them, i.e., the optimal $\mathbf{g}$ is given as:
$\mathbf{g}\in\arg\min_{\mathbf{g}}\mathrm{D}[\mathbb{P}_{\mathbf{X}(\mathbf{g}),\mathbf{M}}|\mathbb{P}_{\mathbf{X}(\mathbf{g})}\otimes\mathbb{P}_{\mathbf{M}}],$ (4)
@@
where $\mathrm{D}[\mathbb{P}|\mathbb{Q}]$ stands for the Kullback-Leibler (KL) divergence of probability distributions $\mathbb{Q}$ from $\mathbb{P}$. However, the mutual information is not directly computable as we do not have access to the true distribution $\mathbb{P}_{\mathbf{X}(\mathbf{g}),\mathbf{M}}$ and $\mathbb{P}_{\mathbf{X}(\mathbf{g})}\otimes\mathbb{P}_{\mathbf{M}}$. A potential solution is to transform this optimization problem into a bi-level min-max adversarial optimization problem: Fixing $\mathbf{g}$, it is possible to estimate the mutual information using samples from $\mathbb{P}_{\mathbf{X}(\mathbf{g}),\mathbf{M}}$ and $\mathbb{P}_{\mathbf{X}(\mathbf{g})}\otimes\mathbb{P}_{\mathbf{M}}$ using Mutual Information Neural Estimation (MINE) (Belghazi et al., 2018) which maximizes the Donsker-Varadhan lowerbound (Donsker and Varadhan, 1975) to approximate the KL divergence. After that, we can minimize the estimated mutual information with respect to $\mathbf{g}$. Repeat until convergence. If one solves both optimization problems one gradient step at a time, this algorithm is the classic adversarial training that GAN is known for.
@@
However, contrary to the suggestion in (Liu et al., 2024), the mutual information in (4) cannot be minimized simply by directly applying forward or reverse KL Wasserstein Gradient Flow (WGF). While WGF can minimize a KL divergence $\mathrm{D}[\mathbb{P}\|\mathbb{Q}]$ when the distribution to be optimized is either $\mathbb{P}$ or $\mathbb{Q}$, the distribution of interest in our setting ($\mathbb{P}_{\mathbf{X}(\mathbf{g})}$) appears simultaneously in both $\mathbb{P}$ and $\mathbb{Q}$. As a result, neither the forward nor reverse KL formulation of WGF is directly applicable. Therefore, while the WGF procedure proposed in (Liu et al., 2024) is conceptually appealing, it does not correctly minimize the mutual information objective as formulated in (4).
@@
Motivated by the limitation of both adversarial training and WGF, we propose a sequential imputation algorithm that reduces the mutual information between $\mathbf{X}^{(t)}$ and $\mathbf{M}$, called Mutual Information Reducing Iterations (MIRI). Let us denote the joint probability of the pair $(\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M})$ as $\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}$ and the product measure of $\mathbf{X}^{(t-1)}$ and $\mathbf{M}$ as $\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}$. For $t=1\dots T$, the algorithm performs the following steps:
@@
    $\mathbf{g}_{t}\in\arg\min_{\mathbf{g}}\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$.
@@
The difference between the first step and (4) is that we replace $\mathbb{P}_{\mathbf{X}(\mathbf{g})}\otimes\mathbb{P}_{\mathbf{M}}$ in (4) with $\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}$, the product measure of the previous iteration. We can also confirm that MIRI does indeed reduce the mutual information:
@@
See Appendix A for the proof. In Section 5.1, we show that the GAIN algorithm can be viewed as an approximate implementation of the above MIRI algorithm.
@@
Let $p_{\mathbf{g}}(\mathbf{x},\mathbf{m})$ be the density of $(\mathbf{X}^{(t)}(\mathbf{g}_{t}),\mathbf{M})$ and $q(\mathbf{x})$ be the density of $\mathbf{X}^{(t-1)}$. $\mathbf{g}\in\arg\min_{\mathbf{g}}\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$ if and only if
@@
The proof can be found in the Appendix D. This result shows that the optimal $\mathbf{g}$ should sample from the conditional distribution of the missing components given the non-missing components according to the marginal distribution of $\mathbf{X}^{(t-1)}$, the imputed data in the previous iteration. (5) inspires us to construct a flow-based generative model with a terminal distribution $q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})$.
As we discuss in Section 5.2 and 5.3, (5) is also the key design principle behind many classic and contemporary data imputation algorithms.
@@ 3.2 Imputation by Rectified Flow
In this section, we show how to construct an optimal imputer for MIRI using an imputation ODE obtained via Rectified Flow training. Let us zoom in on Step 1 of the MIRI algorithm at iteration $t$.
@@
We can define an imputation process by using the optimal velocity field $\mathbf{v}^{*}$. Given a partially observed vector where the missing entries are padded with zeros, i.e., $[\boldsymbol{0},\mathbf{x}_{\mathbf{m}}]$, the imputation process dynamics are governed by the following ODE:
@@
Moreover, we can show that, $\mathbf{g}^{*}$ is an optimal imputer to the MIRI algorithm that we introduced in Section 3.1.
@@
$\mathbf{g}^{*}\in\arg\min_{\mathbf{g}}\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right].$
The main technical challenge of the proof is to show that the ODE (7) can indeed establish a “marginal-preserving” process from $\mathbb{P}_{\mathbf{X}^{(t-1)}|\mathbf{M}}$ to $\mathbb{P}_{\mathbf{X}^{(t-1)}}$. This is different from the classic proof since both our training objective and ODE are different from the regular rectified flow. See Appendix B for the proof. With an imputer $\mathbf{g}^{*}$, we can proceed to the second step of the MIRI algorithm.
Algorithm 1 MIRI with Rectified Flow (Single Imputation)
[…]
1:  $(\mathbf{X}^{(0)},\mathbf{M})\leftarrow\verb|initial_impute|(\tilde{\mathbf{X}},\mathbf{M})$.
2:  for $t=1$ to $T$ do
3:     Initialize a neural network model $\mathbf{v}$.
4:     for $n=1$ to $N$ do
8:        $\mathbf{v}_{t}\leftarrow\verb|SGD_update|\left(\nabla_{\mathbf{v}}\sum_{j}\left\|\mathbf{Y}^{(j)}-\mathbf{v}(\mathbf{X}_{\tau_{j}}^{(j)},\mathbf{M}_{0}^{(j)}\odot\mathbf{X}_{0}^{(j)},\mathbf{M}_{0}^{(j)}\odot\mathbf{X}_{1}^{(j)},\mathbf{M}_{0}^{(j)},\tau_{j})\right\|^{2}\right)$
9:     end for
10:     $\mathbf{X}^{(t)}\leftarrow\mathbf{M}\odot\mathbf{X}^{(t-1)}+(1-\mathbf{M})\odot\verb*|ODE_Solver|((1-\mathbf{M})\odot\mathbf{v},\mathbf{X}^{(t-1)},\mathbf{M})$
11:  end for
12:  return  $\{\mathbf{X}^{(T)}_{i}\}$
@@ 3.3 Initial Condition
To run the imputation process, we require at least a sample of $\mathbf{X}_{0,1-\mathbf{m}}\,\big{|}\,\{\mathbf{X}_{0,\mathbf{m}}=\mathbf{x}_{\mathbf{m}},\,\mathbf{M}=\mathbf{m}\}$. By the definition of $\mathbf{X}_{0}$, this is equivalent to sampling from $\mathbf{X}^{(t-1)}_{1-\mathbf{m}}|\{\mathbf{X}^{(t-1)}_{\mathbf{m}}=\mathbf{x}_{\mathbf{m}},\mathbf{M}=\mathbf{m}\}$. This poses no difficulty, since at each iteration $t$, we maintain the access to the joint sample $(\mathbf{X}^{(t-1)},\mathbf{M})$, from which we can extract the relevant conditional sample by slicing. However, in the first iteration, we need samples from $\mathbf{X}^{(0)}_{1-\mathbf{m}}|\{\mathbf{X}^{(0)}_{\mathbf{m}}=\mathbf{x}_{\mathbf{m}},\mathbf{M}=\mathbf{m}\}$ to kickstart the MIRI algorithm.
One can sample from any distribution as those initial samples and in our experiment, we simply draw independent samples from the normal $N(0,1)$ or uniform $U(0,1)$ to fill out each missing component. Alternatively, one can use another imputation algorithm to suggest initial samples.
4 Practical Implementation
@@
The imputation ODE is solved using the Euler method (see Appendix H). However, other methods may also be used. The Euler method with 100 steps generally provides good performance.
@@ 5 MIRI and Other Imputation Algorithms
In this section, we show that how MIRI is related to other existing imputation algorithms.
@@ 5.1 GAIN
Let $q(\mathbf{x},\mathbf{m})$ be the density function of the joint distribution of $(\mathbf{X}^{(t-1)},\mathbf{M})$. Suppose the discriminator is well-trained in the first step in the GAIN algorithm, i.e., $f(m_{j},\mathbf{x},\mathbf{m}_{-j})\approx q(m_{j}|\mathbf{x},\mathbf{m}_{-j})$, and the generator training in the second step (without the reconstruction error) is
$\mathbf{g}_{t}=\arg\min_{\mathbf{g}}\mathbb{E}\left[\sum_{j}\log q(M_{j}|\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M}_{-j})\right]=\arg\min_{\mathbf{g}}\mathbb{E}\left[\log\prod_{j}q(M_{j}|\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M}_{-j}).\right]$
Using the pseudo-likelihood approximation (Besag, 1975), $\prod_{j}q(M_{j}|\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M}_{-j})\approx q(\mathbf{M}|\mathbf{X}_{\mathbf{g}}^{(t)})$, thus:
@@
$\mathbb{E}\left[\log\frac{q(\mathbf{M},\mathbf{X}_{\mathbf{g}}^{(t)})}{q(\mathbf{X}_{\mathbf{g}}^{(t)})p(\mathbf{M})}\right]\leq\mathbb{E}\left[\log\frac{p_{\mathbf{g}}(\mathbf{M},\mathbf{X}_{\mathbf{g}}^{(t)})}{q(\mathbf{X}_{\mathbf{g}}^{(t)})p(\mathbf{M})}\right]=\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right],$
thus, we can see that GAIN learns its imputer $\mathbf{g}_{t}$ by approximately minimizing a lower bound of $\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$.
There are two approximations made in the above argument: the pseudo-likelihood approximation and the Gibbs’ inequality. First, since GAIN only performs one gradient step update to the imputer $\mathbf{g}_{t}$ at each iteration, the gap between $\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$ and its lower bound may not be large. Second, the pseudo-likelihood approximation can be a good approximation, for example, when the correlation among $M_{j}$ is not too strong. The discriminator training of GAIN resembles a dimension-wise strategy that has been widely used in Ising model estimation (Ravikumar et al., 2010; Meinshausen and Bühlmann, 2006). Note that the optimization in (10) is fundamentally intractable due to the high-dimensional density $q(\mathbf{m}|\mathbf{x}^{(t)}(\mathbf{g}))$. $\mathbf{m}$ is a binary variable defined on $\{0,1\}^{d}$, thus, the density not have a tractable normalizing constant. This further restricts us to the dimension-wise pseudo-likelihood approach in GAIN algorithm.
@@
Recall that Proposition 2 states that minimizing the KL divergence $\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$ is equivalent to choosing a $\mathbf{g}$ such that $p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})=q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})$, where $p_{\mathbf{g}}(\mathbf{x},\mathbf{m})$ is the density of $(\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M})$ and $q(\mathbf{x})$ is the density of $\mathbf{X}^{(t-1)}$. One can simply build an imputer that samples from $q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})$ to impute the missing components. Then, by construction, $p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})=q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}}).$
@@
    Let $\mathbf{g}_{t}$ be a Gibbs sampler for $\mathbb{P}_{\mathbf{X}^{(t-1)}_{1-\mathbf{m}}\mid\mathbf{X}^{(t-1)}_{\mathbf{m}}=\mathbf{x}_{\mathbf{m}}}$ using $\{f_{j}\}$.

### v3

@@ 3.1 Mutual Information Reducing Iterations (MIRI)
We study the problem of training an imputer $\mathbf{g}$ to reduce the dependency between the imputed sample $\mathbf{X}(\mathbf{g})$ and missingness mask $\mathbf{M}$. The approach adopted by (liu24minimizing) is to minimize the mutual information between them, i.e., the optimal $\mathbf{g}$ is given as:
$\mathbf{g}\in\arg\min_{\mathbf{g}}\mathrm{I}(\mathbf{X}(\mathbf{g});\mathbf{M})=\arg\min_{\mathbf{g}}\mathrm{D}[\mathbb{P}_{\mathbf{X}(\mathbf{g}),\mathbf{M}}\|\mathbb{P}_{\mathbf{X}(\mathbf{g})}\otimes\mathbb{P}_{\mathbf{M}}],$ (6)
where $\mathrm{D}[\mathbb{P}\|\mathbb{Q}]$ is the Kullback-Leibler (KL) divergence of probability distributions $\mathbb{Q}$ from $\mathbb{P}$.
@@
However, the mutual information is not directly computable as we do not have access to the true distribution $\mathbb{P}_{\mathbf{X}(\mathbf{g}),\mathbf{M}}$ and $\mathbb{P}_{\mathbf{X}(\mathbf{g})}\otimes\mathbb{P}_{\mathbf{M}}$. A potential solution is to transform this optimization problem into a bi-level min-max adversarial optimization problem: Fixing $\mathbf{g}$, it is possible to estimate the mutual information using samples from $\mathbb{P}_{\mathbf{X}(\mathbf{g}),\mathbf{M}}$ and $\mathbb{P}_{\mathbf{X}(\mathbf{g})}\otimes\mathbb{P}_{\mathbf{M}}$ with Mutual Information Neural Estimation (MINE) (belghazi18a), which maximizes the Donsker-Varadhan lower bound (donsker1975) to approximate the KL divergence. After that, we can minimize the estimated mutual information with respect to $\mathbf{g}$. This process is repeated until convergence. If one solves both optimization problems one gradient step at a time, this algorithm is the classic adversarial training that GAN is known for.
@@
Contrary to the suggestion in (liu24minimizing), the mutual information in (6) cannot be minimized simply by directly applying forward or reverse KL Wasserstein Gradient Flow (WGF). While WGF can minimize a KL divergence $\mathrm{D}[\mathbb{P}\|\mathbb{Q}]$ when the distribution to be optimized is either $\mathbb{P}$ or $\mathbb{Q}$, the distribution of interest in our setting ($\mathbb{P}_{\mathbf{X}(\mathbf{g})}$) appears simultaneously in both $\mathbb{P}$ and $\mathbb{Q}$. As a result, neither the forward nor reverse KL formulation of WGF is directly applicable. Therefore, while the WGF procedure proposed in (liu24minimizing) is conceptually appealing, it does not correctly minimize the mutual information objective as formulated in (6).
@@
Motivated by the limitation of both adversarial training and WGF, we propose a sequential imputation algorithm that reduces the mutual information, called Mutual Information Reducing Iterations (MIRI). Instead of targeting the unknown marginal distribution $\mathbb{P}_{\mathbf{X}(\mathbf{g})}$, we use the imputed data distribution from the previous iteration, $\mathbf{X}^{(t-1)}$, as a stable target. Let us denote the joint probability of the pair $(\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M})$ as $\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}$ and the product measure of $\mathbf{X}^{(t-1)}$ and $\mathbf{M}$ as $\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}$.
For each iteration $t=1\dots T$, the algorithm performs the following steps:
@@
    Find the Optimal Imputer: $\mathbf{g}_{t}\in\arg\min_{\mathbf{g}}\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$.
@@
    Update Data: $\mathbf{X}^{(t)}\stackrel{{\scriptstyle\text{impute}}}{{\leftarrow}}\mathbf{X}^{(t)}(\mathbf{g}_{t}),t\leftarrow t+1$.
We can confirm that MIRI does indeed reduce the mutual information:
@@
See Appendix B for the proof. In Section 4.1, we show that the GAIN algorithm can be viewed as an approximate implementation of the above MIRI algorithm.
@@
Let $p_{\mathbf{g}}(\mathbf{x},\mathbf{m})$ be the density of $(\mathbf{X}^{(t)}(\mathbf{g}_{t}),\mathbf{M})$ and $q(\mathbf{x})$ be the density of $\mathbf{X}^{(t-1)}$. $\mathbf{g}\in\arg\min_{\mathbf{g}}\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$ if and only if
@@
The proof can be found in the Appendix E. This result shows that the optimal $\mathbf{g}$ should sample from the conditional distribution of the missing components given the non-missing components according to the marginal distribution of $\mathbf{X}^{(t-1)}$, the imputed data in the previous iteration. (7) inspires us to construct a flow-based generative model with a target distribution $q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})$.
As we discuss in Section 4.2 and 4.3, (7) is also the key design principle behind many classic and contemporary data imputation algorithms.
@@ 3.2 Imputation by Rectified Flow
In this section, we show how to construct an optimal imputer for MIRI using an imputation ODE obtained via rectified flow training. The optimality condition of Proposition 2 requires an imputer to sample from target distribution $q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})$. The standard diffusion models are restrictive, as they are constrained to a fixed Gaussian reference. We therefore use rectified flow to transport samples from the current imputation (reference) to the target distribution.
We now zoom in on Step 1 of the MIRI algorithm at iteration $t$.
@@
We can define an imputation process by using the optimal velocity field $\mathbf{v}^{*}$. Given a partially observed vector where the missing entries are padded with zeros, i.e., $\boldsymbol{0},\mathbf{x}_{\mathbf{m}}]$, the imputation process dynamics are governed by the following ODE:
@@
We can show that, $\mathbf{g}^{*}$ is an optimal imputer to the MIRI algorithm that we introduced in Section 3.1.
@@
$\mathbf{g}^{*}\in\arg\min_{\mathbf{g}}\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right].$ (12)
The main technical challenge of the proof is to show that the ODE (9) can indeed establish a “marginal-preserving” process from $\mathbb{P}_{\mathbf{X}^{(t-1)}|\mathbf{M}}$ to $\mathbb{P}_{\mathbf{X}^{(t-1)}}$. This is different from the classic proof since both our training objective and ODE are different from the regular rectified flow. See Appendix C for the proof. With an imputer $\mathbf{g}^{*}$, we can proceed to the second step of the MIRI algorithm.
@@ 3.3 Initial Condition
To run the imputation process, we require at least a sample of $\mathbf{X}_{0,1-\mathbf{m}}\,\big|\,\{\mathbf{X}_{0,\mathbf{m}}=\mathbf{x}_{\mathbf{m}},\,\mathbf{M}=\mathbf{m}\}$. By the definition of $\mathbf{X}_{0}$, this is equivalent to sampling from $\mathbf{X}^{(t-1)}_{1-\mathbf{m}}|\{\mathbf{X}^{(t-1)}_{\mathbf{m}}=\mathbf{x}_{\mathbf{m}},\mathbf{M}=\mathbf{m}\}$. This poses no difficulty, since at each iteration $t$, we maintain access to the joint sample $(\mathbf{X}^{(t-1)},\mathbf{M})$, from which we can extract the relevant conditional sample by slicing. However, in the first iteration, we need samples from $\mathbf{X}^{(0)}_{1-\mathbf{m}}|\{\mathbf{X}^{(0)}_{\mathbf{m}}=\mathbf{x}_{\mathbf{m}},\mathbf{M}=\mathbf{m}\}$ to kickstart the MIRI algorithm.
One can sample from any distribution for those initial samples and in our experiment, we simply draw independent samples from the normal $N(0,1)$ or uniform $U(0,1)$ to fill out each missing component. Alternatively, one can use another imputation algorithm to suggest initial samples.
3.4 Practical Implementation
@@ 4 MIRI and Other Imputation Algorithms
In this section, we show how MIRI is related to other existing imputation algorithms.
@@ 4.1 GAIN
Let $q(\mathbf{x},\mathbf{m})$ be the density function of the joint distribution of $(\mathbf{X}^{(t-1)},\mathbf{M})$. Suppose the discriminator is well-trained in the first step in the GAIN algorithm, i.e., $f(m_{j},\mathbf{x},\mathbf{m}_{-j})\approx q(m_{j}|\mathbf{x},\mathbf{m}_{-j})$, then the generator training in the second step (without the reconstruction error) is
$\mathbf{g}_{t}=\arg\min_{\mathbf{g}}\mathbb{E}\left[\sum_{j}\log q(M_{j}|\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M}_{-j})\right]=\arg\min_{\mathbf{g}}\mathbb{E}\left[\log\prod_{j}q(M_{j}|\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M}_{-j}).\right]$ (13)
Using the pseudo-likelihood approximation (Besag1975), $\prod_{j}q(M_{j}|\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M}_{-j})\approx q(\mathbf{M}|\mathbf{X}_{\mathbf{g}}^{(t)})$, thus:
@@
$\mathbb{E}\left[\log\frac{q(\mathbf{M},\mathbf{X}_{\mathbf{g}}^{(t)})}{q(\mathbf{X}_{\mathbf{g}}^{(t)})p(\mathbf{M})}\right]\leq\mathbb{E}\left[\log\frac{p_{\mathbf{g}}(\mathbf{M},\mathbf{X}_{\mathbf{g}}^{(t)})}{q(\mathbf{X}_{\mathbf{g}}^{(t)})p(\mathbf{M})}\right]=\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right],$ (15)
thus, we can see that GAIN learns its imputer $\mathbf{g}_{t}$ by approximately minimizing a lower bound of $\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$.
There are two approximations made in the above argument: the pseudo-likelihood approximation and the Gibbs’ inequality. First, since GAIN only performs one gradient step update to the imputer $\mathbf{g}_{t}$ at each iteration, the gap between $\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$ and its lower bound may not be large. Second, the pseudo-likelihood approximation can be a good approximation, for example, when the correlation among $M_{j}$ is not too strong. The discriminator training of GAIN resembles a dimension-wise strategy that has been widely used in Ising model estimation (Ravikumar2010; Meinshausen2006). Note that the optimization in (14) is fundamentally intractable due to the high-dimensional density $q(\mathbf{m}|\mathbf{x}^{(t)}(\mathbf{g}))$. $\mathbf{m}$ is a binary variable defined on $\{0,1\}^{d}$, thus, the density does not have a tractable normalizing constant. This further restricts us to the dimension-wise pseudo-likelihood approach in GAIN algorithm.
@@
Recall that Proposition 2 states that minimizing the KL divergence $\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$ is equivalent to choosing a $\mathbf{g}$ such that $p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})=q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})$, where $p_{\mathbf{g}}(\mathbf{x},\mathbf{m})$ is the density of $(\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M})$ and $q(\mathbf{x})$ is the density of $\mathbf{X}^{(t-1)}$. One can simply build an imputer that samples from $q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})$ to impute the missing components. Then, by construction, $p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})=q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}}).$
@@
    Let $\mathbf{g}_{t}$ be a Gibbs sampler for $q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})$ using $\{f_{j}\}$.